Timothy owns 980 shares of Countess Corp., which is priced at $14.41 per share. The company plans a 1-for-4 reverse stock split. How many shares will Timothy own and what will the share price be after the reverse stock split?
Answer #1

Calculation of number of shares Timothy own:

Number of shares= 980*1/4= 245 shares

So Timothy will own 245 shares

Calculation of share price:

Share value before split= 980*14.41= 14121.8

Share price after stock split= share value before split/number of shares= 14121.8/245= 57.64

Share price after reverse stock spit is $57.64